OpenStudy (anonymous):

A right cone has a volume 100pie and height of 12. whats the total area?

OpenStudy (anonymous):

What's the formula for a cone?

OpenStudy (anonymous):

Using that, how can we derive the radius?

OpenStudy (anonymous):

So, using the two formulas generously donated above.....\[V=100\pi=\frac{\pi r^2(12)}{3}=\implies 5=r \implies A=\pi r \sqrt{5^2+12^2}=65 \pi\]

Directrix (directrix):

The volume of the cone is given by the formula V = (1/3) B h where B is the area of the base. V = (1/3) pi*r^2 * h 100 pi = (1/3) pi * r^2 (12) 100 pi = 4 pi r^2 25 = r^2 5 = r --------- SA = pi * r * slant height + B SA = pi * (5) * (13) + pi * 5^2 SA = 65 pi + 25 pi SA = 90 pi

OpenStudy (anonymous):

Nice work. I didn't check the formula very well, I guess.

OpenStudy (anonymous):

Just to prove I can do it correctly...\[V=100π=\frac{πr^2(12)}3⟹5=r⟹A=πr (\sqrt{5^2+12^2}+r)=90π\]